OpenStudy (anonymous):

Derek’s favorite clothing store is having their annual fall sale. All clothing is 25 percent off. A shirt he really likes has an original price of $28. Derek has $24 in his wallet. He is wondering if he has enough money to buy the shirt now that it is on sale. He knows he must also pay sales tax. The sales tax rate is 8 percent and is calculated on the discounted price. Does Derek have enough money to buy the shirt? Justify your answer. Your explanation should contain answers to the following questions: 1. What is the amount of the discount?

OpenStudy (littlebird):

The first step is calculating the discounted price, which is 28-(28x.25)

OpenStudy (littlebird):

$21

OpenStudy (littlebird):

No you have to add sales tax

OpenStudy (littlebird):

*Now

OpenStudy (littlebird):

By the way, the discount was $7

OpenStudy (littlebird):

Or 25%

OpenStudy (danjs):

Welcome to Open Study!

OpenStudy (danjs):

All clothing is 25 percent off. The original price of $28. 25% = 0.25 $28 - (0.25)* $28 = $28 - $7 = $21 The price of the item after discount of 25% , is 21 dollars. Another way to think of it, You are paying 75% of the cost, or 0.75 of 28 (0.75)*28 = 21 dollars

OpenStudy (danjs):

The sales tax rate is 8 percent and is calculated on the discounted price 8% = 0.08 Need to add on 8% of that 21 dollars at tax. $21 + 0.08*$21 = $21 + $1.68 = $22.68 The total price after discount and tax is $22.68 If He has 24 dollars, he has enough for the item.

OpenStudy (danjs):

Do you see everything, or not get part of it?